# Beyond Prompt-Induced Lies: Investigating LLM Deception on Benign Prompts

### v1

@@ 4.2 Example of Prompts to LLM
The full prompt posed to the LLM is presented in Figure 3(a). This prompt is derived directly from a complete, linked sequence of individuals, presented in Figure 3(b). For linked-list questions, no “broken edge” is introduced; all facts required to form the logical chain are provided. As a result, the model should apply the transitive rule across the entire sequence, making the expected correct answer to the query “Fiona Flores can contact Colin Miller?” a “Yes”.
@@ 3. If not specified in the facts that A can contact B, A cannot contact B
4.3 Evaluating Deception
Evaluating Deceptive Intention.
Based on Definition 3.3 and Equation 1, we operationalize our evaluation using a pair of tasks with comparable difficulty: a linked-list question ($Q_{L}$) and a broken-linked-list Question ($Q_{B}$) of the same size $n$. A consistent performance gap between these two question types, as measured by our score, indicates the presence of a hidden objective. The nature of this implicit goal is revealed by the sign of $\rho(n;\mathcal{M})$:
  • A positive score ($\rho>0$) reveals a model tendency to complete paths, indicating a deceptive intention achieved by fabrication.
  • A negative score ($\rho<0$) reveals a model tendency to break paths, indicating a deceptive intention achieved by concealment.
Evaluating Deceptive Behavior.
5 Experiment
5.1 Experimental Setup
This section details the datasets, models, metrics, hyperparameters, and implementation used in this study.
Data.
Our evaluation dataset consists of questions generated according to the framework in Section 4.3. We generate 1,000 questions for each combination of question category and length, where the number $n$ of individuals is varied across the set $\{2,3,5,10,20,30,40,80\}$. The experiments utilize five distinct categories, though not all experiments require every category.
  • Linked: A standard linked-list question with a “Yes” ground truth.
  • Broken-Repeat: A broken-linked-list question ($b=\lfloor n/k\rfloor$) that is immediately followed by a repetition of the same query, but with the specific fact corresponding to the broken edge now included.
Models.
We evaluate a diverse set of 14 state-of-the-art LLMs, encompassing both closed and open-source releases from leading organizations such as OpenAI, Microsoft, Google, DeepSeek, Alibaba, Meta, and MistralAI. Table 1 provides a comprehensive list of the models used and their attributes. DeepSeek-V3-0324 is presented separately because it fails to answer the question even at $n=2$ (see Appendix A.1).
Table 1: Details of language models evaluated in this study.
Vendor Model Name Version Size Type
OpenAI o4-mini 2025-04-16 Unknown Closed-Source
o3-mini 2025-01-31 Unknown Closed-Source
gpt-4.1 2025-04-14 Unknown Closed-Source
gpt-4.1-mini 2025-04-14 Unknown Closed-Source
gpt-4o 2024-08-06 Unknown Closed-Source
gpt-4o-mini 2024-07-18 Unknown Closed-Source
Microsoft phi-4 2024-05-14 14.7B Open-Source
Google gemma-2-9b-it 2024-08-28 9B Open-Source
DeepSeek DeepSeek-V3-0324 2025-03-24 685B Open-Source
Alibaba Qwen3-235B-A22B 2025-05-21 235B Open-Source
Qwen3-30B-A3B 2025-05-21 30B Open-Source
Qwen2.5-32B-Instruct 2024-09-25 32B Open-Source
Meta Llama-3.1-8b-instruct 2024-07-23 8B Open-Source
MistralAI Mistral-Nemo-Instruct 2024-07-18 12.2B Open-Source
Metrics.
Our evaluation is guided by the two-dimensional framework of deception outlined in Section 4.3. We operationalize this with two primary metrics. To measure deceptive behavior, we calculate the accuracy gap between responses to an initial query and its follow-up, thereby capturing behavioral inconsistency. To measure deceptive intention, we use the Deceptive Intention Score $\rho$, as formally defined in Equation 1. To report the stability of these measurements, we utilize a bootstrapping procedure with 1,000 sampling iterations to estimate the mean and variance for each metric.
Hyperparameters and Implementation.
We access all models through their respective inference APIs. We query proprietary models from OpenAI via the official OpenAI API and use the API services of the Nebius Platform [^2] https://studio.nebius.com for open-source models. To ensure consistency and reproducibility, we standardize all hyperparameters. As our preliminary analysis shows that model temperature has a negligible impact on the results (see Appendix Figure 11), we set it to $1.0$ for all experiments. Similarly, we set the hyperparameter $k=2$. As established in Section 5.7, this choice does not significantly affect the relative performance ranking, allowing us to maintain a consistent protocol while reducing computational costs. Finally, due to computational constraints, we set the maximum difficulty level to $t=80$ for calculating the metrics $\bar{\delta}$ and $\bar{\rho}$.
5.2 Overall Analysis
In this section, we present the overall deceptive behavior score ($\bar{\delta}$) and deceptive intention score ($\bar{\rho}$) for all models, analyzing their evolution with model size and release date (Figure 5). Our analysis yields three key observations.
First, modern LLMs show progress in reducing deception on simpler tasks, but this improvement is not universal. As shown in Figure 5(a), both scores are minimal for several advanced models, such as o3-mini and the Qwen3 series. Figures 5(b) and 5(c) further illustrate a general decreasing trend in deceptive tendencies with more recent model release dates. However, this progress is inconsistent. For instance, gpt-4.1 exhibits worryingly high scores for both metrics, despite being released alongside the more robust o4-mini. It must be noted that even the best-performing models are not entirely honest; as detailed in Appendix A.1, models like o4-mini still show an increased tendency toward deception as question complexity grows.
Second, higher LLM capacity does not always result in better honesty.. A direct comparison between gpt-4o and gpt-4o-mini, or gpt-4.1 and gpt-4o, reveals that the larger, more powerful models do not consistently demonstrate greater honesty (i.e., scores closer to zero). Instead, their behavior sometimes indicates a shift from one type of error, like systematic hallucination, towards another, like random guessing.
Third, LLMs’ performance on both metrics is positively correlated. Figure 5(a) clearly shows that models with a low deceptive behavior score ($\bar{\delta}$) also tend to have a low absolute deceptive intention score ($\bar{\rho}$). This strong correlation across a diverse set of models supports our hypothesis that these two metrics are linked by the underlying latent variable of deception, confirming that behavioral inconsistency and strategic intent often emerge in parallel.
(a) Overall $\bar{\delta}$ and $|\bar{\rho}|$ ($t\in[2,80]$)
(b) $\bar{\delta}$ evolution over time
(c) $|\bar{\rho}|$ evolution over time
Figure 5: Overall analysis of deceptive behavior and intention across models. (a) Shows the overall deceptive behavior score $\bar{\delta}$ and the absolute deceptive intention score $|\bar{\rho}|$. (b) Shows the evolution of $\bar{\delta}$ over time. (c) Shows the evolution of $|\bar{\rho}|$ over time.
(a) Deceptive intention score $\rho(n,\mathcal{M})$
(b) Deceptive behavior score $\delta(n,\mathcal{M})$
Figure 6: Deception evolution as $n$ increases across all evaluated models with 95% confidence intervals. (a) Shows the trend of deceptive intention score, while (b) presents the trend of deceptive beha score.
5.3 Deceptive Intention
5.4 Deceptive Behavior
First, deceptive behavior emerges as question difficulty increases. When the size $n$ of individuals is small, most models exhibit low deceptive behavior scores. However, as $n$ escalates, the deceptive behavior score rises across all models. The point at which deceptive behavior emerges is contingent on the model’s capability. Stronger models, such as o4-mini and Qwen3-235B-A22B, demonstrate deceptive behavior at $n=20$, whereas weaker models like gpt-4.1-mini and gpt-4o-mini show this behavior at $n=5$.
gpt-4.1 0.415 0.036 0.174 0.269
gpt-4.1-mini 0.715 0.275 0.533 0.617
gpt-4o 0.449 0.005 0.174 0.280
gpt-4o-mini 0.379 0.182 0.584 0.470
o3-mini 0.000 0.011 0.104 0.000
o4-mini 0.000 0.000 0.001 0.000
Table 2: Deceptive Behavior Scores on rephrased questions ($n=10$)
5.5 Joint Analysis of Both Metrics
Since deception is co-determined by deceptive intention and deceptive behavior, we study how both metrics changes as $n$ increases, with results grouped by model family in Figure 7, 8, 9, and 10, respectively.
These figures reveal a key observation: the Deceptive Behavior Score ($\delta$) and the absolute Deceptive Intention Score ($|\rho|$) are highly positively correlated. This strong positive correlation, which holds across most models regardless of vendor, size, or reasoning capability, supports our hypothesis that deception is jointly determined by behavioral inconsistency and strategic intention. The only notable exception is o4-mini, where this trend is not significant because the magnitudes of both $\delta$ and $|\rho|$ are negligible, indicating a high degree of honesty within the tested range ($n\leq 80$). Overall, this widespread correlation confirms that deception emerges systematically as problem complexity increases.
(a) GPT-4.1
(b) GPT-4.1-mini
(c) GPT-4o
(d) GPT-4o-mini
Figure 7: Joint analysis of deceptive intention and behavior for GPT model variants with varying $n$
(a) o3-mini
(b) o4-mini
(c) Phi-4
Figure 8: Joint analysis of deceptive intention and behavior for OpenAI advanced models and Microsoft Phi-4 with varying $n$
(a) Qwen2.5-32B
(b) Qwen3-30B
(c) Qwen3-235B
Figure 9: Joint analysis of deceptive intention and behavior for Qwen model family with varying $n$
(a) Gemma-2-9B
(b) Llama-3.1-8B
(c) Mistral-Nemo
Figure 10: Joint analysis of deceptive intention and behavior for open-source models with varying $n$
5.6 Effect of Temperature
(a) GPT-4.1
(b) GPT-4.1-mini
(c) GPT-4o
(d) GPT-4o-mini
5.7 Effect of Initial-Followup Difficulty Ratio
5.8 Evolution of Deception by Model Size
In this section, we analyze how deception evolves with model size by plotting the deceptive behavior score ($\bar{\delta}$) and the absolute deceptive intention score ($|\bar{\rho}|$) against the number of parameters for open-source LLMs. As shown in Figure 13, we observe a clear and concerning trend: both deceptive behavior and deceptive intention scores tend to increase with model size.
(a) Overall $\bar{\delta}$ vs model size
(b) Overall $|\bar{\rho}|$ vs model size
Figure 13: Analysis of deceptive scores across different model sizes. The x-axis shows the number of parameters (in billions) on a logarithmic scale, while the y-axis represents the respective deception scores.
5.9 Analysis of Deception in Chain-of-Thought
@@
5.10 Visualization of Embeddings
This section visualizes how inconsistent responses, which implies deceptive behavior, emerge and cluster within the model’s internal representations as task difficulty ($n$) increases. Figures 17, 18, and 19 display PCA-reduced embeddings, from which we draw two key findings.
@@ 6 Broader Impact
Our findings indicate a tendency for LLMs to be more deceptive when performing more difficult tasks. Although lacking definitive evidence, we suspect this correlation may not be coincidental. LLMs may be more deceptive on difficult tasks precisely because the deception is harder to verify. This possibility warrants significant attention from the AI community. When deploying LLMs for highly challenging tasks (e.g., proving unsolved mathematical theorems or implementing complex software systems), there may be a higher probability that the model will fabricate a specious lemma or conceal an edge case with conditional logic. Regardless of the model’s underlying intention”, which could simply be to generate a more complete-looking answer, this vulnerability must be addressed before deploying LLM-driven systems in critical roles.
@@
The deceptive behaviors observed in this study suggest that current training objectives may inadvertently teach LLMs to “appear correct” rather than to “be correct and honest.” This implies that the learning goal may be excessively utilitarian, prioritizing plausible outputs over factual integrity. We suspect this behavior is deeply rooted in the pre-training objectives rather than being an artifact of post-hoc alignment fine-tuning. This possibility raises fundamental questions and calls for a re-evaluation of the training paradigms for LLMs.
@@
7 Conclusion
Acknowledgement
We are grateful to Yue Bi for her helpful suggestions regarding the experimental design in the psychological perspective.
Appendix A Appendix
A.1 Deceptive Intention and Behavior Analysis Results
A.1.1 Deceptive Intention Analysis Results
Figures 2023 illustrate the Deceptive Intention Score ($\rho$) for all models across question categories of varying difficulty. We observe two key patterns from these results. First, a consistent trend across most models is that the deceptive intention score tends to escalate with question difficulty. Even the best-performing model, o4-mini (Figure 21(b)), exhibits this pattern, suggesting that increased complexity systematically induces a higher propensity for deception. Second, DeepSeek-V3 (Figure 22(a)) presents a notable exception. Contrary to the general trend, it shows an unusually high failure rate on simple questions. This issue persists in our validation on DeepSeek official website  [^3] https://chat.deepseek.com/. We hypothesize that this anomaly stems from the model’s challenges in comprehending English-language questions.
(a) GPT-4.1
(b) GPT-4.1-mini
(c) GPT-4o
(d) GPT-4o-mini
Figure 20: Deceptive Intention Score ($\rho$) analysis with confidence intervals for GPT model variants
(a) o3-mini
(b) o4-mini
(c) Phi-4
Figure 21: Deceptive Intention Score ($\rho$) analysis for OpenAI advanced models and Microsoft Phi-4
(a) DeepSeek-V3
(b) Gemma-2-9B
(c) Llama-3.1-8B
(d) Mistral-Nemo
Figure 22: Deceptive Intention Score ($\rho$) analysis for prominent open-source models
(a) Qwen2.5-32B
(b) Qwen3-30B
(c) Qwen3-235B
Figure 23: Deceptive Intention Score ($\rho$) analysis for the Qwen model family. This comparison reveals how model scale affects deceptive intention.
A.1.2 Deceptive Behavior Score Analysis
Figures 2427 display the Deceptive Behavior Score ($\delta$) for all evaluated models. The results presented here, which quantify the models’ manifested deceptive actions, are broadly consistent with the analysis of the Deceptive Intention Score ($\rho$) in the main paper.
(a) GPT-4.1
(b) GPT-4.1-mini
(c) GPT-4o
(d) GPT-4o-mini
Figure 24: Deceptive Behavior Score ($\delta$) analysis with confidence intervals for GPT model variants.
(a) o3-mini
(b) o4-mini
(c) Phi-4
Figure 25: Deceptive Behavior Score ($\delta$) analysis for OpenAI advanced models and Microsoft Phi-4
(a) DeepSeek-V3
(b) Gemma-2-9B
(c) Llama-3.1-8B
(d) Mistral-Nemo
Figure 26: Deceptive Behavior Score ($\delta$) analysis for prominent open-source models
(a) Qwen2.5-32B
(b) Qwen3-30B
(c) Qwen3-235B
Figure 27: Deceptive Behavior Score ($\delta$) analysis for the Qwen model family

### v2

@@ Appendix E Example of Prompts of CSQ Framework
The full prompt posed to the LLM is presented in Figure 14(a). This prompt is derived directly from a complete, linked sequence of individuals, presented in Figure 14(b). For linked-list questions, all facts required to form the logical chain are provided. As a result, the model should apply the transitive rule across the entire sequence, making the expected correct answer “Yes” to the query “Fiona Flores can contact Colin Miller?”.
@@ 3. If not specified in the facts that A can contact B, A cannot contact B
Appendix F Case Study: Reasoning behind Deception
F.1 Analysis of Deception in Chain-of-Thought
@@
F.2 Visualization of Embeddings
This section visualizes how inconsistent responses, which implies deceptive behavior, emerge and cluster within the model’s internal representations as task difficulty ($n$) increases. Figures 19, 20, and 21 display PCA-reduced embeddings, from which we draw two key findings.
@@ Appendix G Broader Impact
Our findings indicate a tendency for LLMs to be more deceptive when handling more difficult tasks. Despite the lack of definitive evidence, we suspect this correlation may not be coincidental. LLMs may be more deceptive on difficult tasks precisely because the deception is harder to verify. This possibility warrants significant attention from the AI community. When deploying LLMs for highly challenging tasks (e.g., proving unsolved mathematical theorems or implementing complex software systems), there may be a higher probability that the model will fabricate a specious lemma or conceal an edge case with conditional logic. Regardless of the model’s underlying intention, which could simply be to generate a more complete-looking answer, this vulnerability must be addressed before deploying LLM-driven systems in critical roles.
@@
The deceptive behaviors observed in this study suggest that current training objectives may inadvertently teach LLMs to “appear correct” rather than to “be correct and honest.” This implies that the learning goal may be excessively utilitarian, prioritizing plausible outputs over factual integrity. We suspect this behavior is deeply rooted in the pre-training objectives rather than being an artifact of post-hoc preference alignment, which raises fundamental questions and calls for a re-evaluation of the training paradigms for LLMs.
@@
Appendix H Large Language Model Usage
Large language models played an assistive role in the implementation and polishing of this paper. Code implementation was primarily assisted by Claude-4-Sonnet and subsequently reviewed by Gemini-2.5-Pro for correctness. Manuscript polishing was assisted by Gemini-2.5-Pro and GPT-5.